OpenStudy (anonymous):

The slope of a line that is perpendicular to the line y - 3x = 12, is:

OpenStudy (mattfeury):

what is the slope of this line?

OpenStudy (saifoo.khan):

perpendicular line as the slope, the inverse of the original line!

OpenStudy (anonymous):

so..-1/3?

OpenStudy (mattfeury):

i don't think -3 is quite the slope ;)

OpenStudy (saifoo.khan):

NO, look. the original slope is: -3. so not insert this value in the formulae: \[\large m_1 * m_2 = -1\] \[-3 * m_2 = -1 \] \[m_2 = \frac{-1}{-3} \] \[m_2 = \frac{1}{3}\]

OpenStudy (anonymous):

oooh

OpenStudy (mattfeury):

the original slope is 3. you need to put it in slope-intercept form: y = mx + b

OpenStudy (saifoo.khan):

LOl, my bad.. chnge tht -3 into +3!! sorry! so the final answer will be: \[m= \frac{-1}{3}\]

OpenStudy (anonymous):

you can think of that with angles.. m = 3 => tan(x) = 3 (=) x = 71.56º now you just need to add or subtract 90º it's like 71.56+-90 = +-18.43 tan(+-18.43) = +-1/3 personally I think the graphical way is more simple to understand